OpenStudy (anonymous):

Which rational number equals 0.6 repeating

OpenStudy (anonymous):

let x=0.6666... -(1) 10 x=6.6666... -(2) (2)-(1) gives 9x=6 \[x=\frac{ 6 }{ 9 }=?\]